# MindGYM: Enhancing Vision-Language Models via Synthetic Self-Challenging Questions

### v1

@@ 4.3 Ablation Studies
To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies.
GPT-based Thinking Quality Scoring.
Datasets Depth Breadth Avg
MindGYM/ Raw 10.2% 19.4% 14.8% $\uparrow$
MindGYM/ LIMO 1.63% 23.8% 12.7% $\uparrow$
MindGYM/ Open-O1 (400) 1.88% 23.7% 12.8% $\uparrow$
MindGYM/ MMEvol-DVQA 8.41% 37.0% 22.7% $\uparrow$
Overall 5.53% $\uparrow$ 26.0% $\uparrow$ 15.8% $\uparrow$
Table 2: The results of win rates in terms of relative improvements over baselines, using GPT-4 as a scorer.
We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer. The protocol assesses: (1) Reasoning Depth via derivation complexity, and (2) Reasoning Breadth by solution diversity. Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.5. Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark. The experimental outcomes documented in Table 2 demonstrate MindGYM’ consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in reasoning depth over the raw model and other competitive baselines.
@@
MindGYM-Text w/o CF 52.20 60.83 56.51 $\downarrow$
@@
MindGYM-OKVQA 53.70 61.09 57.40
MindGYM-OKVQA w/o SC 52.85 61.30 57.08 $\downarrow$
MindGYM-OKVQA Syn-EN 53.14 60.59 56.87 $\downarrow$
MindGYM-OKVQA with RB 53.33 60.60 56.97 $\downarrow$
MindGYM-SciQA 53.11 61.46 57.29
MindGYM-SciQA w/o SC 53.20 60.63 56.91 $\downarrow$
MindGYM-SciQA Syn-EN 52.25 60.88 56.56 $\downarrow$
MindGYM-SciQA with RB 53.46 60.27 56.87 $\downarrow$
Table 3: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB), complete results detailed in Appendix C.
Ablation of Stream of Consciousness.
We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same curriculum fine-tuning method for training and evaluation. The experimental results, as shown in Table 3, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reductions ranging from 0.32 to 1.48 points. This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of reasoning and overall performance of the model.
Curriculum Fine-tuning Order.
To verify the effectiveness of our Curriculum Fine-tuning module, we mix and shuffle the first three datasets ($\mathcal{D}^{1}_{\text{train}}$, $\mathcal{D}^{2}_{\text{train}}$, $\mathcal{D}^{3}_{\text{train}}$) and train the model on text data, subsequently using $\mathcal{D}^{4}_{\text{train}}$ to constrain the answers’ generation pattern. This experimental setup allows us to systematically examine whether disrupting the curriculum order impacts model performance. The experimental results shown in Table 3 under the w/o CF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the Curriculum Fine-tuning order on model performance.
@@
Based on the selected Qwen series [4] model, we utilize Chinese as the language for data synthesis. Additionally, we explore the effects of using English for synthesis, as shown in Table 3 under the Syn-EN condition. The results indicate varying degrees of performance degradation when using English, with a decrease of up to 1.31 points in the MindGYM-Text setting. Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it more beneficial for the model’s reasoning learning.
@@
To verify the effectiveness of allowing the model to autonomously select relationship categories, we balance the number of entries generated for each category, totaling 400 entries. The experimental results shown in Table 3 under the with RB condition indicate that enforcing category balance leads to a decline up to 1.33 in overall model performance. This finding suggests that the autonomous selection of relationships by the model reflects its cognitive understanding.
@@ 5 Conclusion & Future Works
This work introduces MindGYM, a framework that enhances vision-language reasoning via synthetic self-challenging questions. By self-challenging multi-step reasoning data and employing structured curriculum training, MindGYM achieves strong performance gains with small data size and reduced computational costs, outperforming prior methods in reasoning depth and breadth.
Future directions include extending cognitive relationships to dynamic visual scenarios [46]; exploring the application of MindGYM in specific domains like medicine [27] and finance; developing adaptive scoring operators [6] for question complexity; and integrating agents [13] for data correctness verification in generalized RL environments.
Appendix A Implementation of Experiments
A.1 Utilized Seed Visual Datasets
For the multimodal domain, we leverage widely-adopted scientific and reasoning datasets as visual source material for question synthesis, specifically including ScienceQA and OK-VQA.
A.2 Benchmarks
@@
A.3 Training and Evaluation Details
@@
    • LIMO [43]: Challenges conventional assumptions in mathematical reasoning by demonstrating that models achieve superior performance with smaller quantities of high-quality training data. We reformat its ”question+solution+answer” structure into standard instruction pairs for training consistency.
@@
Appendix B Prompts
B.1 Stream of Consciousness
@@
B.2 Prompts for Background Generation
@@
B.3 Prompts for Seed Single-Hop Question Synthesis
@@
B.4 Prompts for Challenging Multi-Hop QA Synthesis
@@
B.5 Prompts for GPT-based Scorer
@@ 3. Comparative analysis: use bullet points to list the main improvements/degradations
Please return a JSON that strictly follows this format: ”accuracy_raw”: 0-2, ”accuracy_finetuned”: 0-2, ”depth_raw”: 0-3, ”depth_finetuned”: 0-3, ”comparison”: [”Point 1”, ”Point 2”]
Appendix C Complete Results including Ablation Study
We present the full experimental results in Table 4, comprehensively validating the effectiveness of our method compared to other baselines, as well as the robustness of our approach itself (ablation studies introduced in Section 4.3).
@@ Appendix C Complete Results including Ablation Study
Table 4: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB).
Appendix D Alternative VLMs Data Synthesis through MindGYM
As demonstrated in Listing 1, our analysis using InternVL2.5-8B [5] reveals its generated outputs tend to exhibit chaotic patterns. We hypothesize this stems from dual factors: (i) constrained input token limits restricting contextual comprehension, and (ii) inherent deficiencies in cognitive reasoning capabilities. These limitations necessitated its exclusion from our comparative VLM experiments.
Listing 1: Data Sample of MindGYM on InternVL2.5-8B.
1{
2 "background_document": "Background document for multi-hop reasoning.",
3 "reasoning_category": "Mathematical Reasoning",
4 "sub_questions": [
5 "What is the optimal pricing strategy considering cost inflation?",
6 "How will fixed and variable costs vary with seasonal demand?",
7 "What are the impacts of geopolitical changes on supply chain?",
8 "What role do regulatory changes play in pricing?",
9 "How will budget allocation differ among different product lines?"
10 ],
11 "relationship_category": "Comparison",
12 "multihop_question": "What steps need to be taken to optimize pricing strategies taking into account rising costs, shifting regulations, seasonal fluctuations, and differing product revenues?",
13 "multihop_answer": "To maximize profit under increased costs, companies should reassess fixed and variable expenses, implement strategic cost savings wherever feasible, monitor regulatory developments closely, review existing budgets, allocate resources accordingly, update pricing schedules, communicate revisions internally consistently, measure effectivity quarterly.",
[…]
15}

### v2

@@ 4.4 Ablation studies
To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies on Qwen2.5-VL-7Bqwen25vl .
Ablation of Stream of Consciousness.
Table 4: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of our fine-tuning (w/o OF) steps and Relation Balanced (with RB), complete results detailed in Appendix D.
@@
MindGYM-Text w/o OF 52.20 60.83 56.51 $\downarrow$
@@
We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same fine-tuning paradigm for training and evaluation. The experimental results, as shown in Table 4, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reduction of 1.48 points. This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of thinking and overall performance of the model.
The Usage of Dataset.
To verify the effectiveness of our fine-tuning module, we mix and shuffle the first three phase and train the LLM on text data, subsequently using the data of autonomous solving phase to constrain the answers’ generation pattern. This experimental setup allows us to systematically examine whether disrupting the fine-tuning order impacts model performance. The experimental results shown in Table 4 under the w/o OF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the our utilized order on the synthesis data.
@@
Based on the selected Qwen qwen25vl , we utilize Chinese as the language for data synthesis. Additionally, we explore the effects of using English for synthesis, as shown in Table 4 under the Syn-EN condition. The results indicate the performance degradation when using English, with a decrease of 1.31 points. Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it beneficial for the model’s thinking learning.
@@
To verify the effectiveness of allowing the model to autonomously select relationship categories, we balance the number of entries generated for each category, totaling 400 entries. The experimental results shown in Table 4 under the with RB condition indicate that enforcing category balance leads to a decline of 1.33 in overall performance. This finding suggests that the autonomous selection of relationships by LLM reflects its cognitive understanding.
@@ 5 Conclusion & Future Works
We present MindGYM, a cognitively guided framework for synthesizing self-challenging vision-language data. By injecting high-level reasoning signals and composing multi-hop tasks, MindGYM enables models to acquire structured thinking with minimal data and computation. Our synthetic data achieves higher quality and lower variance, leading to strong gains across reasoning benchmarks.
Future directions include extending cognitive relationships to dynamic visual scenarios zhou2024humanvbenchexploringhumancentricvideo ; exploring the application of MindGYM in specific domains like medicine MURTAZA2023100546 and finance; developing adaptive scoring operators dj2 for question complexity; and integrating agents as for data correctness verification in generalized RL environments.
Appendix
Our appendix is organized as follows:
  • Sec. A: The details of experimental setup, including the benchmarks (Sec. A.1) and the training & evaluation details (Sec. A.2).
  • Sec. B: The implementation details of MindGYM, including the description of meta-topics (Sec. B.1) and all prompts utilized in this paper (Sec. B.2).
  • Sec. C: Our exploration on multimodal data synthesis with MindGYM, including our proposed method on multimodal data (Sec. C.1), the experimental setup (Sec. C.2), experimental results (Sec. C.3) and the comprehensive analysis of the synthesis data (Sec. C.4).
  • Sec. D: The ablation stuides, including the full experimental results.
Appendix A Detailed Experimental Setup
A.1 Benchmarks
@@
A.2 Training and Evaluation Details
@@
    • LIMO [39]: Challenges conventional assumptions in mathematical reasoning by demonstrating that models achieve superior performance with smaller quantities of high-quality training data. We reformat its "question+solution+answer" structure into standard instruction pairs for training consistency.
@@
Appendix B Implementation Details of MindGYM
B.1 Detailed Description of Meta-topics
Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
  • Historical Events: Analyzing past events and their consequences.
  • Economic Principles: Studying resource allocation and valuation.
  • Psychological Insights: Exploring human behavior and cognition.
For purely textual contexts, we define three canonical combination types:
  • Comparison: Identifying contrasts or similarities across entities or events;
  • Bridging: Linking information across distant knowledge pieces via intermediate inference;
  • Temporal: Reasoning across sequential or causal timelines.
These types serve as explicit cues in $P_{\text{task}}^{\text{multi-hop}}$, guiding $\pi$ to adopt different reasoning schemata.
B.2 Utilized Prompts
B.2.1 Stream of Consciousness
@@
B.2.2 Prompts for Background Generation
@@
B.2.3 Prompts for Seed Single-Hop Question Synthesis
@@
B.2.4 Prompts for Challenging Multi-Hop QA Synthesis
@@
B.2.5 Prompts for GPT-based Scorer
@@ 3. Comparative analysis: use bullet points to list the main improvements/degradations
Please return a JSON that strictly follows this format: "accuracy_raw": 0-2, "accuracy_finetuned": 0-2, "depth_raw": 0-3, "depth_finetuned": 0-3, "comparison": ["Point 1", "Point 2"]
Appendix C Multimodal Data Synthesis Exploration
C.1 Methodologies
Anchoring Image Sources.
Multimodal Cognitive Combination Types.
C.2 Experimental Setup
While the main body of our work focuses exclusively on synthetic textual data, we also conduct preliminary explorations into multimodal data synthesis to demonstrate the extensibility of our approach. These experiments employ the same set of vision-language models—Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-32B-Instruct, InternVL-8B, and InternVL-38B—for generating synthetic image-question-answer triplets.
For the multimodal domain, we leverage widely-adopted scientific and reasoning datasets as visual source material for question synthesis, specifically including ScienceQA and OK-VQA. These datasets serve as visual grounding primitives, enabling the models to synthesize multimodal questions that align with real-world visual contexts.
During training with image-text data, we update only the Aligner-layers, keeping both the ViT and LLM layers frozen to preserve pretrained representations and ensure stable adaptation. These results are not the main focus of the current study but demonstrate the extensibility of our method to multimodal contexts.
Table 5: Performance of different models on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training. Different from Table 1, this table is about multimodal data synthesized by MindGYM.
Text Eval Multimodal Eval
Model Series Dataset GSM8K Math GPQA Text-Avg MMStar MathVista MathVision MM-Avg Avg
[…]
Openo1-sft (400) 84.31 69.00 29.70 61.00 63.87 69.70 23.66 53.08 57.04
Openo1-sft (4k) 77.94 59.00 28.19 55.04 58.93 61.40 17.43 45.92 50.48
[…]
MMEvol-SciQA (106) 83.40 65.60 31.83 60.28 63.93 69.50 23.36 52.26 56.27
MMEvol-DvQA (4k) 83.85 65.80 33.08 60.91 62.93 67.40 24.01 51.45 56.18
[…]
Openo1-sft (400) 95.38 81.20 50.51 75.70 68.80 73.40 36.18 59.46 67.58
[…]
MMEvol-SciQA (106) 95.60 81.00 42.42 73.01 69.00 73.80 36.84 59.88 66.44
MMEvol-DvQA (4k) 92.72 80.60 38.38 70.57 69.20 72.60 39.14 60.31 65.44
[…]
LIMO (817) 89.16 76.40 42.42 69.33 69.00 73.40 32.57 58.32 63.82
[…]
MMEvol-DvQA (4k) 88.55 74.60 40.40 67.85 68.47 71.60 32.24 57.44 62.65
[…]
LIMO (817) 89.16 72.60 47.47 69.74 72.40 72.80 36.51 60.57 65.16
MMEvol-SciQA (106) 89.46 75.00 46.46 70.31 72.40 73.20 38.49 61.36 65.83
MMEvol-DvQA (4k)} 89.92 74.00 46.46 70.13 70.93 71.70 30.92 57.85 63.99
MindGYM-OKVQA(ours) 88.93 75.60 48.99 71.17 72.53 71.80 36.18 60.17 65.67
[…]
C.3 Experimental Results
Cross-modal Transfer.
Generalizability & Robustness.
C.4 Data Analysis
Table 6: Data-Juicer analysis results across different models. Different from Table 2, this table is about multimodal data synthesized by MindGYM.
Model Dataset
quality
-mean
quality
-std
action dependency token length
baseline Openo1-sft (400) 0.96 0.091 8.05 2.00 77 274
Openo1-sft (4k) 0.70 0.10 8.78 2.02 83 284
LIMO (817) 0.87 0.10 7.71 2.02 101 322
MMEvol-SciQA (106) 0.91 0.082 2.31 1.85 13.73 67
MMEvol-DvQA (4k) 0.76 0.11 2.66 1.95 14.82 69
Qwen2.5-VL-7B MindGYM-OKVQA 0.93 0.079 9.04 2.30 122 96
MindGYM-SciQA 0.90 0.099 10.23 2.27 144 117
Qwen2.5-VL-32B MindGYM-OKVQA 0.97 0.044 10.94 2.25 168 134
MindGYM-SciQA 0.98 0.031 15.44 2.16 233 195
InternVL-8B MindGYM-OKVQA 0.96 0.058 9.62 2.32 141 110
MindGYM-SciQA 0.96 0.045 10.64 2.26 148 118
InternVL-38B MindGYM-OKVQA 0.74 0.069 7.19 2.28 111 91
MindGYM-SciQA 0.97 0.039 9.26 2.23 131 119
We conduct a comparative analysis across baselines and our MindGYM-generated datasets to assess the impact of cognitively guided data on reasoning quality and complexity. As shown in Table 6, MindGYM datasets (OKVQA and SciQA) consistently yield higher reasoning quality scores, with Qwen2.5-VL-32B achieving the best performance (0.98 quality mean with 0.031 std on MindGYM-SciQA), significantly outperforming standard instruction-tuning sets such as Openo1-sft (0.70) and LIMO (0.87). In addition, MindGYM prompts models to generate substantially longer and more structured reasoning traces, as evidenced by the increase in action steps (up to 15.44) and dependency depth (avg. >2.2). This indicates that our cognitively annotated data effectively stimulates deeper reasoning behaviors without sacrificing generation stability. Notably, while token and input length grow with model capacity, the reasoning quality remains robust, suggesting that MindGYM scales well with model size and supports more abstract reasoning. In contrast, traditional instruction datasets lead to flatter, shorter outputs with limited cognitive structure. These results demonstrate the value of our data-centric approach in enhancing model reasoning ability and interpretability.
Appendix D Comprehensive Ablation Studies
We present the complete version of Table 4 in Table 7, which demonstrates the comprehensive ablation study results of our method. In addition to conducting ablation experiments on the text-based variant of our approach (as described in Appendix C), we also performed an ablation study on the multimodal data synthesis methodology. As shown in Table 7, the performance degradation observed in the OKVQA and SciQA ablation experiments compared to the full implementation of our method provides compelling evidence for two key aspects: 1) the effectiveness of our proposed framework, and 2) its cross-modal generalization capability across different data modalities. This empirical validation underscores the importance of our multimodal synthesis strategy in maintaining model performance across diverse vision-language tasks.
@@ Appendix D Comprehensive Ablation Studies
Table 7: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of fine-tuning (w/o OF) steps and Relation Balanced (with RB).